Click Here to receive Service Notes automatically via email.
| Date | Description |
|---|---|
| 08/29/2008 | The storm Thursday evening took down electrical power to many locations in the Kirksville area, including several Truman buildings. During the electrical outages some Truman IT services were unavailable. When electrical power was restored all IT services were also restored. |
| 08/29/2008 | Friday evening at 9 PM during the regularly scheduled IT maintenance window system updates will be installed on the Blackboard system. Blackboard services will be unavailable while the updates are being installed. We estimate Blackboard services will be unavailable for an hour. |
| 08/13/2008 | IT Services successfully completed the installation of new collaboration tools for Blackboard tonight. This upgrade caused no downtime. The first workshop featuring the new tools is scheduled for Truman week (8/22/08 - 9:30am) Sign up at: http://tctl.truman.edu/register/ |
| 08/12/2008 | ITS will be working to install some additional tools for use in Blackboard overnight tonight (8/12/08) between 11pm and 1am. Little if any down time is anticipated. |
| 08/08/2008 | During tonight's (08/08/08) regularly scheduled ITS maintenance window, an update to the Blackboard login page referencing browser compatibility is planned. The update will be applied at 5:30pm and we are anticipating 0-10 minutes of downtime. |
| 07/25/2008 | Network services have been restored in Pershing Building. |
| 07/25/2008 | Network services have been restored to Magruder Hall. |
| 07/25/2008 | Network services to the old Grim Smith Hospital building are currently unavailable due to telephone wiring problems caused by the heavy rain. |
| 07/25/2008 | Barnett telephone repairs are scheduled to be completed today (Friday). |
| 07/15/2008 | There was a brief faculty / staff email outage at ~ 11 AM this morning due to a server error. The problem has been resolved and email services restored. |
| 06/25/2008 | Last night around 11:23pm, the Banner database server suffered a system crash. The nightly backup and system maintenance scripts recovered and brought ADMIN backup though its listener and a couple of the support services (such as posting) were not brought backup. As of about 8:25am, Banner ADMIN services were restored. We will work with IBM support on identifying the source of the system crash.
Banner HARRY failed to clone properly last night due to the crashed state of Banner ADMIN. Banner HARRY will not be available today. Banner HARRY should successfully clone with tonights regular backups and system maintenance and be available tomorrow. For those of you working on special projects, Banner BESS is recovered and should be available for use. |
| 06/26/2008 | CableOne has fixed the Internet connection at the University Farm. Truman network services are now available at the University Farm. |
| 06/18/2008 | We experienced a brief Internet outage around 10:35 am. Our monitors showed the Internet as unavailable for ~ 8 minutes. We contacted MOREnet and the problem was on the state network. It has been corrected. |
| 06/11/2008 | Blackboard services will not be available today from 5p - 5:15pm to address problems with the collaboration / chat tools. |
| 06/07/2008 | ITS will be performing server upgrades on Saturday 6/7/2008 from 8am - Noon. Service interruptions will be concentrated within the first hour of work, but will be possible during the entire period. Websites impacted by the maintenance are as follows:
www.truman.edu |
| 06/02/2008 | Truman State Universities Internet service provider, MOREnet, is experiencing congestion on their Internet links in Kansas City and St. Louis. Until the issue is resolved there may be delays in Internet connectivity. |
| 05/30/2008 | The site www.truman.edu was unavailable from approximately 11:45 p.m. on 5/29/2008 until 8:30 a.m. on 5/30/2008 due to an internal server error. All services have been restored, the cause of the outage is being investigated. |
| 05/26/2008 | Our Internet connection is back up. The MOREnet sub-contractor has repaired the damaged equipment in Kansas City. |
| 05/23/2008 | The problem with the primary network file server has been resolved.
You may need to log off and log back on to reconnect to network file services. |
| 05/19/2008 | A number of Truman websites were corrupted this morning due to a successful hacking attempt. Functionality is now restored and we are working to find and fix the security hole. |
| 05/02/2008 | Electrical power is restored. All IT services are back to normal. ----- Electrical power failed again in the data center this morning. Many IT services are currently unavailable. The electricians are on-site and are working to restore power as soon as possible. ----- |
| 04/21/2008 | Sunday morning at 6:30 AM electrical power will be turned off in Violette Hall to work on an electrical problem. Network services will be unavailable to much of campus during this time. Physical Plant estimates it will take several hours to resolve the electrical issue.
A memo from Physical Plant with more information was emailed to all faculty, staff and students on Friday. In general network service to buildings connected to Violette Hall will be unavailable during the power outage. Network service to buildings connected to McClain Hall should stay up while Violette Hall is down. The following web page shows which campus buildings are connected to Violette and which are connected to McClain. https://its.truman.edu/networkservices/images/Network%20Logical%20View.jpg This information is also available from the ITS web site under Network and Accounts, Network Design, Truman Network Diagram. |
| 03/20/2008 | The problem with the intermittent slow Internet has been resolved. |
| 03/18/2008 | The issue with outgoing email has been resolved. The system causing the problem has been removed from the campus network. |
| 03/18/2008 | Banner and therefore all Banner Services was down this morning. A Banner process hung during the nightly backup process. We cleared the problem this more. Banner was down from approximately 4am to 8:25am. |